OpenStudy (anonymous):

an you help me now with these |x + 2| + 16 = 1 i need help :( |2x + 6| − 4 = 20 nd help solving

OpenStudy (anonymous):

Are these two different questions?

OpenStudy (jdoe0001):

\(\large { |x + 2| + 16 = 1\implies |x + 2| =1-16\\\quad\\\implies \begin{cases} +(x+2)=-15\implies x=?\\ -(x+2)=-15\implies x=? \end{cases} }\)

OpenStudy (jdoe0001):

\(\large { |2x + 6| -4 = 20\implies |2x + 6|=20+4 \\ \quad \\ \begin{cases} +(2x+6)=24\implies &x=?\\ -(2x+6)=24\implies &x=? \end{cases} }\)

OpenStudy (anonymous):

i dont know what to do after that